the equation for a market demand curve is Qd = 15 – 0.2P and the equation for a market supply curve is Qs = ̶ 1 + 0.6P, what is the consumer surplus?


1
Expert's answer
2021-01-19T07:22:17-0500

Let's plot demand and supply curves in Excel:



Here, orange curve is the market supply curve, the blue curve is the market demand curve. The point at the intersection of these two curves called the market equilibrium. We can see from the graph that the equilibrium price equals "P_E=20" and equilibrium quantity equals "Q_E=11".

Then, we can find the consumer surplus from the area of the upper triangle:


"CS=\dfrac{1}{2}\cdot Base\cdot Height,""CS=\dfrac{1}{2}\cdot (75-20)\cdot 11=302.5"

Answer:

"CS=302.5"